# Radar-Guided Polynomial Fitting for Metric Depth Estimation

### v1

@@ 4 Experiments
Baseline Methods. We consider five recent radar-camera depth estimation baselines: RadarNet [45], Lin [26], SparseBeatsDense [24], GET-UP [49], and RadarCam-Depth [25]. Notably, RadarCam-Depth also leverages radar points to rescale initial scaleless MDE predictions.
Datasets. We evaluate all methods on the nuScenes, ZJU-4DRadarCam (ZJU), and View-of-Delft (VoD) datasets (see Supp. Mat for more details) using the MAE and RMSE metrics with maximum evaluation distances of 50, 70, and 80 meters, following established conventions in the literature.
@@ 4.1 Main Results
PolyRad reduces MAE by 12.1% and RMSE by 10.5% on nuScenes, 46.8% and 62.4% on ZJU, and 31.8% and 38.5% on VoD, achieving SOTA results for all datasets (see Tab. 1).
@@
We qualitatively compare PolyRad against all baseline methods on the challenging nuScenes dataset, providing a visual demonstration of its improved depth estimation performance (see Fig. 3). PolyRad more accurately predicts metric scale and exhibits fewer global misalignments. Furthermore, it is able to extrapolate depth predictions of object surfaces from learned photometric priors, an ability inherited from the backbone MDE model (UniDepth) trained on large-scale datasets. Our polynomial fitting approach then non-uniformly corrects depth discrepancies across the scene, leading to significant performance improvement over initial MDE predictions. Introducing multiple degrees of freedom enables PolyRad to correct misalignments of objects relative to each other, improving accuracy by correcting consistent over- or under-estimations in local regions. For instance, the raw MDE prediction for Image C incorrectly places the bus stop roof at around the same depth as the more distant tree branch, as highlighted on the right. Additionally, MDE does not accurately infer the boundary between the curb and asphalt, again placing both at erroneously similar depths. With a learned polynomial transformation, PolyRad corrects these misalignments with depth-dependent adjustments, accurately separating the curb from the asphalt and placing the bus stop roof and tree branch at their correct depths.
@@
While RadarCam-Depth also utilizes an MDE backbone, its complex processing of raw MDE predictions can be seen to distort initially correct depth structure. In Image C, the tree trunk and attached fern are incorrectly placed at different depths despite the initial MDE predictions correctly positioning them at similar depths. Furthermore, overfitting to specific regions within a scene—particularly those with denser ground truth—may reduce overall error relative to the MDE model but can lead to unintended geometric artifacts. This can result in the omission of entire structures in the predicted depth map, such as the highlighted building on the left of Image A and the construction crane arms in Image B. GET-UP also struggles to accurately infer depth structure, resulting in the omission of objects across all three samples: the highlighted building in Image A, the crane arms in Image B, and foreground tree branches in Image C.
@@
DPT [41] 5188.2 6884.5
Depth Anything [56] 2404.9 4851.1
Depth Pro [4] 3835.0 6600.3
UniDepth [37] 2129.8 4887.7
@@
PolyRad w/ DPT 1525.6 3745.0
PolyRad w/ Depth Anything 1515.1 3719.4
PolyRad w/ Depth Pro 1627.7 4143.9
PolyRad w/ UniDepth 1454.1 3488.9
Table 2: MDE backbone comparative studies on nuScenes.
4.2 Comparative Studies
MDE Models. We evaluate PolyRad using four monocular depth estimation (MDE) backbones: DPT [41], Depth Anything [56], Depth Pro [4], and UniDepth [37]. DPT and Depth Anything infer scaleless inverse depth, and while Depth Pro and UniDepth output metric depth, their raw predictions exhibit significant deviation from metric ground truth (see Tab. 2 and Supp. Mat), and thus we process them the same way as scaleless depth.
Tab. 2 compares raw MDE performance to their performance when used as backbones for RadarCam-Depth and PolyRad. Raw MDE predictions show large errors due to scale ambiguity, while RadarCam-Depth provides moderate improvements. In contrast, PolyRad consistently reduces error across all backbones, demonstrating the effectiveness of polynomial fitting. For DPT and Depth Anything, outputs are first inverted when used as a backbone, and are further median-scaled for the reported raw results. Among all configurations, PolyRad w/ UniDepth achieves the best performance, improving over raw UniDepth predictions by 30.2% and over RadarCam-Depth w/ UniDepth by 20.8%.
Polynomial Degree. As polynomial degree—selected as a hyperparameter—increases, so does the number of inflection points, allowing PolyRad to correct misalignments between the many scene elements in Fig. 4, including cars, trucks, trees, buildings, and traffic lights. This is reflected in the successively improved error maps. Tab. 3 quantifies the performance gains achieved with higher-degree polynomials. At degree 10, performance degrades slightly, potentially due to excessive flexibility resulting in detrimental oscillations.
Ablations. We evaluate the impact of key components of PolyRad in Tab. 4. Replacing learnable prototypes with self-attention on the radar point features degrades performance, demonstrating that prototypes capture relevant patterns within radar point configurations. Removing cross-modality fusion results in the largest performance drop, highlighting the necessity of leveraging correspondences between MDE and radar features. Removing the monotonicity loss term also reduces performance, suggesting that depth ordering regularization is crucial for stable polynomial fitting.
@@ 4.2 Comparative Studies
Table 3: Sensitivity study of polynomial degree hyperparameter.
@@
no ablations 1454.1 3488.9 1500.1 3951.8
Table 4: Ablation studies of architecture and loss components.
5 Discussion
Limitations. PolyRad requires tuning the polynomial degree as a hyperparameter, which may vary across datasets. Additionally, while our novel first-derivative regularization mitigates oscillatory behavior from high-degree polynomials, further constraints could enhance stability.
Summary. We are the first to formulate radar-camera depth estimation as a scene-fitting problem, leveraging high-degree polynomials to transform MDE predictions into accurate metric depth. Our principled and efficient approach demonstrates that a fundamental insight—shifting from simple scale-and-shift to flexible, radar-guided polynomial transformations—can outperform more complex systems.

### v2

@@ 4 Experiments
Baseline Methods. We consider five recent radar-camera depth estimation baselines: RadarNet (Singh et al., 2023), SparseBeatsDense (Li et al., 2024b), GET-UP (Sun et al., 2025), RadarCam-Depth (Li et al., 2024a), and TacoDepth (Wang et al., 2025). Notably, both RadarCam-Depth and TacoDepth also take scaleless MDE predictions as input, together with radar points, to predict metric depth.
Datasets. We evaluate all methods on the nuScenes, ZJU-4DRadarCam (ZJU), and View-of-Delft (VoD) datasets (see Sec. C for more details) using the MAE and RMSE metrics with maximum evaluation distances of 50, 70, and 80 meters, following established conventions in the literature.
@@ 4.1 Main Results
Compared to baselines, POLAR reduces MAE by 4.4% and RMSE by 3.7% on nuScenes, 38.5% and 57.5% on ZJU, and 31.8% and 38.5% on VoD, achieving SOTA results for all datasets (Tab. 1).
@@
We qualitatively compare POLAR against baseline methods on the challenging nuScenes dataset, providing a visual demonstration of its improved depth estimation performance (Fig. 4). POLAR more accurately predicts metric scale and exhibits fewer global misalignments. Furthermore, it is able to extrapolate depth predictions of object surfaces from learned photometric priors, an ability inherited from the backbone MDE model (e.g., UniDepth) trained on large-scale datasets. Our polynomial fitting approach then non-uniformly corrects depth discrepancies across the scene, leading to significant performance improvement over initial MDE predictions. Introducing multiple degrees of freedom enables POLAR to correct misalignments of objects relative to each other, improving accuracy by correcting consistent over- or under-estimations in local regions. For instance, the raw MDE prediction for Image C incorrectly places the bus stop roof at around the same depth as the more distant tree branches, as highlighted on the right. Additionally, MDE does not accurately infer the boundary between the curb and asphalt, again placing both at erroneously similar depths. With a learned polynomial transformation, POLAR corrects these misalignments with depth-dependent adjustments, accurately separating the curb from the asphalt and placing the bus stop roof and tree branch at their correct depths.
@@
While RadarCam-Depth also utilizes an MDE backbone, its complex processing of raw MDE predictions can be seen to distort initially correct depth structure. As shown in the top-left highlighted region of Image C, the tree trunk and attached fern are incorrectly placed at different depths despite the initial MDE predictions correctly positioning them at similar depths. Furthermore, overfitting to specific regions within a scene—particularly those with denser ground truth—may reduce overall error relative to the MDE model but can lead to unintended geometric artifacts. This can result in the omission of entire structures in the predicted depth map, such as the highlighted building on the left of Image A and the construction crane arms in Image B. GET-UP also struggles to accurately infer depth structure, resulting in the omission of objects across all three samples: the highlighted building in Image A, the crane arms in Image B, and foreground tree branches in Image C.
As the polynomial degree—selected as a hyperparameter—increases, the maximum possible number of inflection points also grows, giving POLAR greater expressive power to correct misalignments between the many scene elements in Fig. 3, including cars, trucks, trees, buildings, and traffic lights. This is reflected in the successively improved error maps. See Sec. A for quantitative evaluations.
4.2 Computational Efficiency
Table 2: Training times (minutes per epoch) on nuScenes using an NVIDIA A6000 GPU.
Method Train Time / Epoch
Lin (Lin et al., 2020) 89.25
RadarNet (Singh et al., 2023) 101.50
SparseBeatsDense (Li et al., 2024b) 63.96
GET-UP (Sun et al., 2025) 249.57
RadarCam-Depth (Li et al., 2024a) 86.38
POLAR (Ours) 33.16
Table 3: Inference time in milliseconds (ms) and computational cost (GFLOPs) on nuScenes.
Method Inference Time GFLOPs
Lin (Lin et al., 2020) 129.96 550.43
SparseBeatsDense (Li et al., 2024b) 97.47 532.74
GET-UP (Sun et al., 2025) 445.45 630.99
RadarCam-Depth (Li et al., 2024a) 315.64 619.02
TacoDepth (Wang et al., 2025) 29.30 139.87
POLAR (Ours) 24.81 89.70
POLAR achieves state-of-the-art training time, inference time (ms), and computational overhead (GFLOPs) compared to all baseline methods, while also achieving state-of-the-art accuracy. Tab. 3 shows that POLAR has the lowest training time per epoch among all methods, requiring 33.16 minutes per epoch on nuScenes. This efficiency stems from our streamlined design, which avoids multi-stage processing and explicit radar-camera association learning, both of which contribute to longer training times in methods such as RadarCam-Depth and GET-UP.
Tab. 3 highlights POLAR’s state-of-the-art inference speed, requiring just 24.81 ms per frame—a reduction of 15.3% compared to the previous state-of-the-art TacoDepth and 92.1% compared to RadarCam-Depth. This corresponds to 40.3 fps, enabling real-time depth perception. Furthermore, POLAR achieves this inference speed with a lower computational overhead of 89.70 GFLOPs—a 39.5% reduction relative to TacoDepth and an 85.5% reduction relative to RadarCam-Depth. Taken together, these results establish POLAR as not only the most accurate radar-camera depth estimation method, but also the most practical for real-time deployment where latency is critical.
5 Discussion
Limitations. POLAR requires tuning the polynomial degree as a hyperparameter (Sec. A), which may vary across datasets. Additionally, while our novel first-derivative regularization mitigates oscillatory behavior from high-degree polynomials, further constraints could enhance stability.
Summary. We are the first to formulate radar-camera depth estimation as a scene-fitting problem, leveraging high-degree polynomials to transform MDE predictions into accurate metric depth. Our principled and efficient approach demonstrates that a fundamental insight—shifting from affine scale-and-shift to flexible polynomial transformations—outperforms computationally heavier methods.
Appendix
Appendix A Comparative Studies
Table 4: MDE backbone comparative studies on nuScenes.
@@ Appendix A Comparative Studies
DPT (Ranftl et al., 2021) 5188.2 6884.5
Depth Anything (Yang et al., 2024) 2404.9 4851.1
Depth Pro (Bochkovskii et al., 2024) 3835.0 6600.3
UniDepth (Piccinelli et al., 2025) 2129.8 4887.7
@@
POLAR w/ DPT 1525.6 3745.0
POLAR w/ Depth Anything 1515.1 3719.4
POLAR w/ Depth Pro 1627.7 4143.9
POLAR w/ UniDepth 1407.8 3193.5
MDE Models. We evaluate POLAR using four monocular depth estimation (MDE) backbones: DPT (Ranftl et al., 2021), Depth Anything (Yang et al., 2024), Depth Pro (Bochkovskii et al., 2024), and UniDepth (Piccinelli et al., 2025). DPT and Depth Anything infer scaleless inverse depth, and while Depth Pro and UniDepth output metric depth, their raw predictions exhibit significant deviation from metric ground truth (see raw MDE performance in Tabs. 41010), and thus we process them the same way as scaleless depth.
Tab. 4 compares raw MDE performance to their performance when used as backbones for RadarCam-Depth and POLAR. Raw MDE predictions show large errors due to scale ambiguity, while RadarCam-Depth provides moderate improvements. In contrast, POLAR consistently reduces error across all backbones, demonstrating the effectiveness of polynomial fitting. For DPT and Depth Anything, outputs are first inverted when used as a backbone, and are further median-scaled for the reported raw results. Among all configurations, POLAR w/ UniDepth achieves the best performance, improving over raw UniDepth predictions by 51.2% and over RadarCam-Depth w/ UniDepth by 37.8%.
Table 5: Sensitivity study of polynomial degree, selected as a hyperparameter.
@@
Table 6: Ablation studies of architecture and loss components.
@@
pos. embeddings 1454.1 3488.9 1500.1 3951.8
no ablations 1407.8 3193.5 629.6 1171.3
Polynomial Degree. Tab. 6 quantifies the performance gains achieved with higher-degree polynomials, with degree 8 outperforming lower degrees. At degree 10, performance degrades slightly, potentially due to excessive flexibility resulting in detrimental oscillations.
Figure 5: To quantify absolute improvement in our qualitative results, we provide the colorbars that were used in all examples in Figs. 346, and 7.
Ablations. We evaluate the impact of key components of POLAR in Tab. 6. Replacing learnable prototypes with self-attention on the radar point features degrades performance, demonstrating that prototypes capture relevant patterns within radar point configurations. Removing cross-modality fusion results in the largest performance drop, highlighting the necessity of leveraging correspondences between MDE and radar features. Removing the monotonicity loss term also reduces performance, suggesting that depth ordering regularization is crucial for stable polynomial fitting.
Figure 6: Additional qualitative comparison on VoD. POLAR more accurately predicts metric scale, and exhibits fewer global misalignments and geometric artifacts in comparison to other methods. Polynomial fitting, allowing POLAR to non-uniformly scale distinct local regions, results in notable improvements in metric depth estimation over initial MDE predictions. This is apparent in all MDE error maps: though object depth geometry is reasonable, actual metric depth values tend to be inaccurate. GET-UP, while marginally improving metric depth estimation relative to MDE, presents substantial errors in depth structure. For instance, the presence of geometric artifacts or omission of objects or regions entirely can be seen through the following highlighted examples: loss of motorcycle body and visor in Image A, front portion of biker (right) and bike wheel (far-left) in Image B, top of motorcyclist in Image C, and street sign in Image D. RadarCam-Depth (RC-D) similarly exhibits geometric artifacts, present in the sign pole, motorcycle, and fence post in Image A, biker and bicycle wheel in Image B, roof overhang and motorcyclist in Image C, and fence posts and pedestrians in Image D. In contrast, POLAR extrapolates correct structure from photometric priors while also demonstrating clear improvements in metric depth fidelity compared to all other baselines.
Appendix B More on Computational Efficiency
Incremental increase in computation for each additional polynomial degree. Adding a $k+1\text{-th}$ polynomial term requires minimal computational overhead. For an MDE prediction of shape $(H,W)$, the additional cost consists of three components: predicting another coefficient via the linear layer of ($1\times 64\times 2$ FLOPs), computing the $(k+1)\text{-th}$ power from pre-computed $k\text{-th}$ exponentiation ($H\times W$ multiplications), and incorporating this term into the final depth prediction ($2\times H\times W$ operations for multiplication by the coefficient and addition). For nuScenes, where each image has shape $(900,1600)$, each additional polynomial term incurs an additional computational cost of $0.0043$ GFLOPs, which is a $0.0048\%$ increase.
Appendix C Dataset Details
[…]
Appendix D Full nuScenes Benchmark
Table 7: Full quantitative results ($m$) on the nuScenes benchmark.
nuScenes
Distance Method MAE RMSE
50m NLSPN (Park et al., 2020) 2790.0 5813.4
BPNet (Tang et al., 2024) 2407.0 4438.0
RC-PDA (Long et al., 2021) 2225.0 4156.5
RC-PDA-HG (Long et al., 2021) 2210.0 4234.0
BTS (Lee et al., 2019) 1937.0 3885.0
DORN (Lo & Vandewalle, 2021) 1926.6 4124.8
RadarNet (Singh et al., 2023) 1727.7 3746.8
CaFNet (Sun et al., 2024) 1674.0 3674.0
Lin (Lin et al., 2020) 1598.2 3790.1
SparseBeatsDense (Li et al., 2024b) 1524.5 3567.3
RadarCam-Depth (Li et al., 2024a) 1286.1 2964.3
GET-UP (Sun et al., 2025) 1241.0 2857.0
TacoDepth (Wang et al., 2025) 1046.8 2487.5
POLAR (Ours) 1014.4 2475.7
70m NLSPN (Park et al., 2020) 3140.0 6580.6
RC-PDA (Long et al., 2021) 3326.1 6700.6
RC-PDA-HG (Long et al., 2021) 3485.6 7002.9
BTS (Lee et al., 2019) 2346.0 4811.0
DORN (Lo & Vandewalle, 2021) 2170.0 4532.0
RadarNet (Singh et al., 2023) 2073.2 4590.7
CaFNet (Sun et al., 2024) 2010.0 4493.0
Lin (Lin et al., 2020) 1897.8 4558.7
SparseBeatsDense (Li et al., 2024b) 1822.9 4303.6
RadarCam-Depth (Li et al., 2024a) 1587.9 3662.5
GET-UP (Sun et al., 2025) 1541.0 3657.0
TacoDepth (Wang et al., 2025) 1347.1 3152.8
POLAR (Ours) 1286.1 2947.3
80m NLSPN (Park et al., 2020) 3257.7 6872.4
RC-PDA (Long et al., 2021) 3721.0 7632.0
RC-PDA-HG (Long et al., 2021) 3664.0 7775.0
AdaBins (Bhat et al., 2021) 3541.0 5885.0
P3Depth (Patil et al., 2022) 3130.0 5838.0
LapDepth (Song et al., 2021) 2544.0 5151.0
PnP (Wang et al., 2018) 2496.0 5578.0
BTS (Lee et al., 2019) 2467.0 5125.0
DORN (Lo & Vandewalle, 2021) 2432.0 5304.0
RadarNet (Singh et al., 2023) 2179.3 4898.7
CaFNet (Sun et al., 2024) 2109.0 4765.0
Lin (Lin et al., 2020) 1988.4 4841.1
SparseBeatsDense (Li et al., 2024b) 1927.0 4609.6
RadarCam-Depth (Li et al., 2024a) 1689.7 3948.0
GET-UP (Sun et al., 2025) 1632.0 3932.0
TacoDepth (Wang et al., 2025) 1492.4 3324.8
POLAR (Ours) 1407.8 3193.5
We present the full set of quantitative results on the nuScenes dataset in Tab. 7. This table includes additional baseline methods that were omitted from the main text for brevity, providing a comprehensive comparison of POLAR against all known existing radar-camera depth estimation methods. As shown, POLAR consistently outperforms all competing methods across all maximum distance thresholds (50m, 70m, and 80m), achieving the lowest mean absolute error (MAE) and root mean squared error (RMSE).
Appendix E Comparison to Depth Completion
One potential idea for radar-camera depth estimation is to apply lidar-camera depth completion methods designed to densify sparse depth maps using surrounding context. One such method, BPNet (Tang et al., 2024) achieves state-of-the-art performance on the KITTI depth completion benchmark by leveraging bilateral propagation. However, when applied to radar-camera depth estimation, BPNet performs poorly, as shown in Tab. 7. POLAR outperforms BPNet by 57.9% in MAE and 44.2% in RMSE on nuScenes, highlighting the limitations of directly applying lidar depth completion methods to radar data. The key reason for this underperformance lies in the fundamental differences between lidar and radar point clouds. Unlike lidar, radar measurements are orders of magnitude sparser and significantly noisier due to factors such as limited antenna elements, elevation ambiguity (see Fig. 7), and lower range resolution (Singh et al., 2023). Lidar depth completion methods assume a relatively dense and structured input (Xia et al., 2023), leveraging local spatial continuity to propagate depth estimates effectively. In contrast, radar points are too sparse for such methods to infer meaningful local depth relationships, leading to poor depth reconstruction when attempting direct densification.
Table 8: Error metrics for depth estimation. These metrics compute the error between predicted depth $\hat{d}(x)$ and ground truth depth $d(x)$.
Metric Definition
MAE $\downarrow$ $\frac{1}{|\Omega|}\sum_{x\in\Omega}|\hat{d}(x)-d(x)|$
RMSE $\downarrow$ $\Big(\tfrac{1}{|\Omega|}\sum_{x\in\Omega}|\hat{d}(x)-d(x)|^{2}\Big)^{1/2}$
In addition, we compare against Non-Local Spatial Propagation Network (NLSPN) (Park et al., 2020), which achieves near state-of-the-art performance on the KITTI depth completion benchmark by refining sparse lidar depth with an iterative non-local spatial propagation procedure. NLSPN predicts an initial depth map along with pixel-wise confidences, then refines it by estimating non-local neighbors and their corresponding affinities to selectively propagate depth information. Unlike other approaches that rely on fixed local neighbors, NLSPN adaptively determines relevant non-local neighbors, improving depth completion accuracy, especially near depth boundaries. However, again, when applied to radar-camera depth estimation, NLSPN performs poorly, as shown in Tab. 7. POLAR outperforms NLSPN by 59.8% in MAE and 55.4% in RMSE on nuScenes.
Instead of relying on depth completion-style densification, POLAR directly learns a transformation function from radar to metric depth by leveraging polynomial fitting. Our method avoids the pitfalls of propagating unreliable local depth information by refining MDE predictions with learned polynomial coefficients, enabling flexible, scene-adaptive depth corrections that effectively capture object relationships and global scene structure.
Figure 7: nuScenes dataset visualization. The elevation ambiguity of radar points results in erroneous projection onto the image plane that makes it challenging for depth completion methods to infer dense depth. In contrast, lidar points yield a denser, image-aligned projection, which is why accurate 3D scene reconstruction with depth completion methods is possible.
Appendix F Evaluation Metrics and Implementation Details
The evaluation metrics used in our study, Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE), are formulated in Tab. 8. Lower values equal better performance for both MAE and RMSE. Unless specified otherwise, all reported values are in millimeters (mm). We train for 60 epochs using a cosine decay learning rate scheduler with learning rate of $5\times 10^{-5}$, and use weighting terms $\lambda_{L_{1}}=1.0,\lambda_{L_{2}}=0.4,\lambda_{mono}=0.25$ for our loss function.
Appendix G Additional Comparisons
Full MDE Comparisons. For ZJU (see Tab. 10), among all configurations, POLAR w/ UniDepth achieves the best performance, improving over raw UniDepth predictions by 61.1% and over RadarCam-Depth w/ UniDepth by 54.2%. For VoD (see Tab. 10), POLAR w/ UniDepth achieves the best performance in MAE, improving over raw UniDepth predictions by 42.4% and over RadarCam-Depth w/ UniDepth by 32.7%, while POLAR w/ Depth Anything achieves the best performance in RMSE, improving over RadarCam-Depth w/ Depth Anything by 37.6% and over raw Depth Anything predictions (inverted and median scaled) by 10.4%.
Table 9: MDE backbone comparative studies on ZJU-4DRadarCam (ZJU).
ZJU
Method MAE RMSE
DPT (Ranftl et al., 2021) 1885.3 3326.1
Depth Anything (Yang et al., 2024) 1943.2 3469.3
Depth Pro (Bochkovskii et al., 2024) 1680.2 3144.9
UniDepth (Piccinelli et al., 2025) 1533.0 3188.4
RadarCam-Depth w/ DPT 1183.5 3229.0
RadarCam-Depth w/ Depth Anything 1724.4 3661.3
RadarCam-Depth w/ Depth Pro 1490.6 3429.5
RadarCam-Depth w/ UniDepth 1152.5 3168.6
POLAR w/ DPT 707.1 1216.9
POLAR w/ Depth Anything 657.2 1225.4
POLAR w/ Depth Pro 640.3 1174.8
POLAR w/ UniDepth 629.6 1171.3
Table 10: MDE backbone comparative studies on View-of-Delft.
View-of-Delft
Method MAE RMSE
DPT (Ranftl et al., 2021) 4117.9 5498.9
Depth Anything (Yang et al., 2024) 3270.5 4411.9
Depth Pro (Bochkovskii et al., 2024) 3275.9 5936.7
UniDepth (Piccinelli et al., 2025) 2605.6 5691.0
RadarCam-Depth w/ DPT 4013.5 5911.9
RadarCam-Depth w/ Depth Anything 3103.6 6328.7
RadarCam-Depth w/ Depth Pro 2843.4 6082.0
RadarCam-Depth w/ UniDepth 2227.4 5385.8
POLAR w/ DPT 1891.4 4252.6
POLAR w/ Depth Anything 1770.5 3951.8
POLAR w/ Depth Pro 1520.2 3987.2
POLAR w/ UniDepth 1500.1 3960.5
Leveraging Radar. Tab. 13 shows that replacing radar points with learnable, dataset-specific points worsens MAE by 28.0% and RMSE by 29.9%, demonstrating that we indeed leverage the radar inputs effectively. As additional evidence, Tab. 12 shows that our method, evaluated zero-shot cross-dataset, achieves comparable or better performance than baselines trained on the target datasets. Tab. 12 shows we are more robust to reduced radar point density at inference, i.e., less performance degradation than the baseline method RadarCam-Depth.
Table 11: Zero-shot generalization.
$\rightarrowZJU nuScenes\rightarrowVoD$
Method MAE RMSE MAE RMSE
GET-UP (zero-shot) 3845.2 8469.7 4809.1 8653.9
RadarCam-Depth (zero-shot) 5435.9 9785.8 7521.5 9194.8
GET-UP (trained) 1699.7 3882.6 2917.3 6145.1
RadarCam-Depth (trained) 1183.5 3229.0 2227.4 5385.8
Ours (zero-shot) 1147.9 3109.5 2256.2 4744.2
Table 12: Reduced radar point density.
RadarCam-Depth Ours
% radar kept / removed MAE RMSE MAE RMSE
25% kept / 75% removed 7969.4 10831.5 2416.4 4836.6
50% kept / 50% removed 4819.8 7077.7 1816.8 3945.7
75% kept / 25% removed 2537.3 4247.4 1575.2 3611.8
100% kept / 0% removed 1689.7 3948.0 1407.8 3193.5
Table 13: Additional comparisons of POLAR vs. learnable dataset-specific points in place of radar points, and regression baselines.
MAE RMSE
Learnable Points 1860.9 4207.1
Isotonic Reg. 2895.2 4340.5
Cubic Hermite 2131.3 4588.0
PCHIP 1809.6 4054.7
LoRA 2030.6 4493.7
ViT-Adapter 1859.6 4280.0
Our Performance 1407.8 3193.5
Adapters. Tab. 13 shows that recent adapter-based finetuning methods LoRA (Hu et al., 2022) and ViT-Adapter (Chen et al., 2022), even with a post-hoc linear fit to projected radar points, do not outperform us.
Regression Baselines. Isotonic regression and monotone spline fitting methods are natural baselines. Tab. 13 shows these methods for regressing projected radar points on MDE predictions do not outperform us. We hypothesize that this is due to noise in radar points that can be mitigated by learning (Sec. 3.2).
Our learned polynomial fit may, in principle, introduce unwanted inversions of initially correct MDE predictions. POLAR successfully mitigates this effect through the proposed novel first-derivative regularization term (see Sec. 3.4 and Eqs. 67), which effectively constrains such inversions. To quantify this, we compute Kendall’s $\tau$ coefficient between predicted and ground-truth depths. Our method achieves the highest monotonicity ($\tau=0.969$) over regression baselines Isotonic Regression ($\tau=0.871$), PCHIP ($\tau=0.758$), and Cubic Hermite Spline ($\tau=0.736$). The raw MDE predictions do exhibit monotonicity with respect to ground truth ($\tau=0.957$), but our polynomial transformation increases it, indicating that we correct unwanted inversions. To assess statistical significance, we compute Kendall’s $\tau$ over 30 bootstrap samples for both our method and the raw MDE predictions. A two-sample t-test reveals a statistically significant difference in mean monotonicity ($p=0.012$).
Appendix H Proof: Limitations of Global Scale and Shift
As further theoretical justification, we prove by construction that an affine scale-and-shift transformation is insufficient to fit MDE predictions to ground truth.
Proposition 1.
There exist infinitely many sets of $k\geq 3$ (MDE prediction $\hat{d}$, ground truth $d$) pairs such that no global scale $\alpha$ and shift $\beta$ satisfy $d=\alpha\hat{d}+\beta$ for all $k$ pairs simultaneously.
Proof by Construction.
Consider the following three (MDE prediction $\hat{d}$, ground truth $d$) pairs from the nuScenes dataset, specifically from the image shown in Fig. 3:
$(\hat{d},d)\in\{(5,7),(10,9),(15,44)\}.$
Assume to the contrary that there exist $\alpha,\beta\in\mathbb{R}$ such that
$d=\alpha\hat{d}+\beta$
holds for all pairs.
From the first two pairs, we obtain:
$7=5\alpha+\beta,\quad 9=10\alpha+\beta.$
Subtracting gives $\alpha=0.4$ and $\beta=5$ as the unique solution for these two pairs.
Applying this solution to the third pair yields:
$\alpha\cdot 15+\beta=0.4\cdot 15+5=11,$
which contradicts the required equality with the ground truth value $d=44$, since the residual error equals 4411=3344-11=33 and not zero.
Hence no global scale $\alpha$ and shift $\beta$ exist that can satisfy all three pairs simultaneously. Moreover, scaling each pair by any nonzero constant produces infinitely many distinct $3$-sets of ($\hat{d},d$) pairs for which no $\alpha$ and $\beta$ exists. Then, for any such $3$-set, appending k3k-3 arbitrary pairs yields an infinite family of $k$-sets ($k>3$) that likewise admit no solution.
Corollary 1.
It is therefore a misconception that the scale ambiguity in MDE can be resolved solely by a global scale and shift. In contrast, any smooth relationship between $\hat{d}$ and $d$ can be locally approximated by a polynomial via Taylor expansion, giving our polynomial fitting formulation the theoretical capacity to approximate $d$ as a function of $\hat{d}$ arbitrarily well.